DaddyDun:

What number needs added to complete the square: x^2+18x+?

celticcat:

divide the coefficient of x by 2 then square it:- coefficient of x = + 18 so we have x^2 + 18x + 81 ( = (x + 9)^2 ) the requred number is 81

FruitDuit:

There is a plus sign where the constant is, meaning both factors will have positive signs. The variable with the coefficient of eighteen will mean there is a difference of half eighteen (this is because the two x factors will be added to make that 18x). 18/2=9. Now we plug in the nine and create (x + 9)(x + 9). Then we use F.O.I.L. First: x * x = x^2 Outer: x * 9 = +9x Inner: x * 9 = +9x Last: 9 * 9 = 81 Added together and put in order you have obtained x^2 + 18x + 81. Conclusion: The coefficient is 81.